a yard in the shape of a square meaures 20ft on each side. a triangular area with a height of 6ft and a base of 5ft is dug up for a flower bed. how much yard area is left over?
Found 2 solutions by jim_thompson5910, solver91311:
Answer by jim_thompson5910(35256)

Area of entire yard = 20*20 = 400 square feet

Area of triangle = (5*6)/2 = 30/2 = 15 square feet


Area left over = 400 - 15 = 385 square feet.

Answer by solver91311(24713)



The area of the square is the measure of a side times itself. The area of the triangle is the measure of the base times the measure of the height divided by 2. Subtract the area of the triangle from the area of the square.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism